OpenStudy (anonymous):

Which fraction is between 1 and 2? is it 1/4 1 1/2 2 2/3 6/8

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

\[1/4 = 0.25\] \[1~ 1/2 \implies 1+1/2 = 1.5\] \[2+2/3 \implies 2.6...\] \[6/8 \implies 0.75\] you could do long division if you have trouble with fractions

OpenStudy (astrophysics):

So what is your answer?

OpenStudy (anonymous):

how is long division

OpenStudy (astrophysics):

I mean you can just divide if you have trouble, \[\frac{ 1 }{ 4 }\] for example, you know how to divide right?

OpenStudy (anonymous):

yea

OpenStudy (astrophysics):

Yeah so you can just do that :P if you're not sure what it is

OpenStudy (anonymous):

B. 1 1/2 is the answer